Heat flows through a wooden board 30mm thick at 0.0086 W/cm² when one of it's sides is 20°C warmer than the other. Determine the thermal conductivity of the wood.​


1
Expert's answer
2020-07-06T15:13:27-0400

Since,

"Q=\frac{kA}{L}\Delta T\implies k=\frac{Q}{A}\frac{L}{\Delta T}"

In this case, "\Delta T=20K,\frac{Q}{A}=0.0086W/cm^2=86W/m^2" ,"L=30mm=0.03m"

Thus, on applying the above formula we get,

"k=86\cdot(0.03/20)=0.129WmK^{-1}"